OpenStudy (anonymous):

If R is the total resistance of two resistors connected in parallel with resistances R1 and R2, then R=R1R2/(R1+R2). If R1 is measured to be 16 ohms with a possible percentage error of 3 percent while R2 is measured to be 34 ohms with a possible percentage error of 5 percent, use differentials to estimate the maximum percentage error in the calculation of R.

OpenStudy (jamesj):

\[dR ={ {\partial R} \ \over { \partial R_1}} dR_1 + { {\partial R} \ \over { \partial R_2}} dR_2\] So you need to calculate this and the substitute into the RHS.

OpenStudy (anonymous):

ya i did that but it say wrong answer

OpenStudy (anonymous):

i found dr/dr1 =r^2/(r1+r2)^2

OpenStudy (jamesj):

dR/dR1 (partials) = [ (R1 + R2)R2 - R1R2 ] / (R1 + R2)^2 = R2^2 / (R1 + R2)^2

OpenStudy (anonymous):

ya its the same i got but im getting wrong answer ?

OpenStudy (jamesj):

you wrote down above r^2/(r1 + r2)^2; did you calculate with r2^2 in the numerator?

OpenStudy (anonymous):

R1 =16 R2=34 dr1 0.03 dr2=0.05

OpenStudy (anonymous):

ya i did r2^2

OpenStudy (jamesj):

ah dr1 is not 0.03. It is 3% of R1. So dr1 = 0.03 x 16 = 0.48

OpenStudy (jamesj):

same for dr2

OpenStudy (anonymous):

ok i got it so i ll put dr2=1.7 right?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

do i need to *100 the ansewr to get % error

OpenStudy (jamesj):

the % error of a variable X is dX/X

OpenStudy (anonymous):

u know im getting wrong anserw too

OpenStudy (anonymous):

i got 0.3960

OpenStudy (anonymous):

and it say wrong ?????????????

OpenStudy (jamesj):

that's far too high. There can't be 39.6% error. You've made a calculation error somewhere along the way.

OpenStudy (jamesj):

(dX/X * 100 is the % error, strictly speaking)

OpenStudy (anonymous):

im doing it again and still same number

OpenStudy (anonymous):

0.396

OpenStudy (anonymous):

can u tell me where im wrong ???????????

OpenStudy (jamesj):

I get 7.7%

OpenStudy (jamesj):

coeff R1 = r1^2/(r1+r2)^2 and r1 error = (coeff r1)*(delta r1)

OpenStudy (anonymous):

ya but i didnt get my coeff r1 is ur coeff r2

OpenStudy (anonymous):

and ur answer is wrong too

OpenStudy (jamesj):

Oh, yes, I reveresed it. One sec

OpenStudy (anonymous):

the error is not delta R

OpenStudy (anonymous):

which u got .396 same as mine ?

OpenStudy (jamesj):

No. We want dR/R * 100. That's the percentage error

OpenStudy (anonymous):

so how u got r

OpenStudy (jamesj):

We have the same dR, which is good. But that's not percentage error. That's absolute error.

OpenStudy (jamesj):

I calculated it using the original formula!

OpenStudy (anonymous):

ok which is 10.88

OpenStudy (jamesj):

Yep

OpenStudy (anonymous):

i got 3.639 %

OpenStudy (jamesj):

Yes, as did I: delta R/R

OpenStudy (anonymous):

and it is correct :)

OpenStudy (jamesj):

good

OpenStudy (anonymous):

thanks :)